OpenStudy (anonymous):

Find the product... (2b+4)(2b-4)

OpenStudy (anonymous):

Use: \[\large \color{green}{(x+y)(x-y) = x^2 - y^2}\] Here x = 2b and y = 4..

OpenStudy (anonymous):

ok substitute 2b for x and 4 for y...So 2b^2 - 16??

OpenStudy (anonymous):

You will also apply square to 2: \[(2b)^2 - (4)^2 \implies 4b^2 - 16\]

OpenStudy (anonymous):

got it...so what is the product? Sorry math and I just do not get along